OpenStudy (anonymous):

can someone help me out and find y' of f(x)= sinxln(5x)

OpenStudy (anonymous):

use product rule

myininaya (myininaya):

use product rule!

OpenStudy (saifoo.khan):

something or the other.

OpenStudy (anonymous):

first function * derivative of second + second function* derivative of first

myininaya (myininaya):

\[y'=(sinx)'\ln{5x}+sinx(\ln{5x})'=cosxln{5x}+sinx \frac{5}{5x}\]